The bungee cord has an unstretched length of 20 m.

The student jumps off the bridge.


At the lowest point in the jump when the student is stationary, the extension of the bungee cord is 35m.

The bungee cord behaves like a spring with a spring constant of 40 N / m.


Calculate the energy stored in the stretched bungee cord




1
Expert's answer
2021-03-04T11:53:57-0500

The energy stored in the stretched cord can be found as follows:


"PE_s=\dfrac{1}{2}kx^2=\dfrac{1}{2}\cdot40\ \dfrac{N}{m}\cdot(35\ m)^2=24500\ J."